OpenStudy (anonymous):

If the side of a square are lengthened by 5 cm, the area become 121 cm^2. please help trying to figuar out the length of the original sqare

OpenStudy (lgbasallote):

equation of the area is s^2. So your equation is (s+5)^2 = 121 where s is the original length of the side of the square. Square root both sides, s + 5 = 11. Therefore, s = 6

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

let the side of the square be 'a' given : (a+5)^2 =121 =>a+5= 11 => a=6